subtract. simplify by collecting like radical terms if possible. 7√32-7√2

Found 2 solutions by JBarnum, mathmandotcom:
Answer by JBarnum(2146)
Answer by mathmandotcom(31)

7*sqrt32-7*sqrt2; Sqrt32=sqrt(16)sqrt(2)=4sqrt(2) by substitution we get
7*4sqrt(2)-7sqrt(2)
28sqrt(2)-7sqrt(2)=21sqrt(2) answer
21sqrt(2).......I hope this helps